# LLMs Can Compensate for Deficiencies in Visual Representations

### v1

@@ Appendix A Dataset Filtering
In order to maintain annotation consistency and visual clarity, we specifically focus on the subset of animal classes within the Pascal Panoptic Parts (Pascal-PP) dataset (de Geus et al., 2021; Meletis et al., 2020) dataset as described in Section 3. The choice is motivated by the relatively uniform and well-defined part structures of animals, as well as their good coverage in the vocabulary of the VLMs we evaluate. To reduce noise and complexity further, we sequentially apply the following filtering criteria:
@@
After filtering, we sample up to 100 images per class, where available. Table 1 provides further details about the distribution of images across different object classes.
@@ Appendix B Additional Results
Figures 10, 11, 12 and 13 presents the layer-wise evolution of per-object part identifiability across different attention knockout settings described in the Section 4.3 for TinyLLaVA 1.1B, BakLLaVA 7B, LLaVA 7B and LLaVA 13B models.
Figures 14, 16 and 15 presents the layer-wise evolution of per-object part identifiability when progressively modifying self-attention across different layers of the CLIP Encoder for BakLLaVA 7B, LLaVA 7B and LLaVA 13B models.
@@ Appendix C Relationship between identifiability and frequency of object-parts
The heatmap in Figure 7 presents the co-occurrence counts between different objects (columns) and their associated parts (rows), whereas the bar plot displays the overall frequency of individual parts irrespective of object association. We can observe that co-occurrence frequencies vary widely across object-part pairs. Common parts such as head, leg, and eye frequently appear with the person class, whereas parts like hoof and wing are more sparsely represented, mainly occurring with specific objects like horse and bird respectively. One might hypothesize that object parts associated with frequently occurring classes – particularly person would exhibit higher identifiability. However, analysis of OPI scores (see Table 3) reveals no clear correlation between co-occurrence frequency and identifiability. In fact, parts with relatively low frequency often achieve high identifiability. Similarly, in Figure 4, parts such as eye in the context of dog and cat, and leg in the context of bird and person are identified accurately despite their lower co-occurrence counts. These findings suggest that the identifiability of object parts in VLMs is not predominantly driven by training data frequency. Instead, VLMs are capable of implicitly learning object-part associations leveraging strong visual or semantic priors, and exhibit effective generalization capabilities, even without explicitly being trained for fine-grained object parts recognition.
@@ Appendix D Relationship between identifiability rate and size of object parts
In this section, we investigate whether the identifiability of object parts is influenced by their relative size. Intuitively, one might expect larger parts to be easier to identify due to their increased visibility and spatial prominence in images. To test this, we consider the CLIP model (with best-performing configuration $l=22$; see Section 4.5) and three LLaVA variants: BakLLaVA 7B, LLaVA 7B, and LLaVA 13B. We begin by binning object parts sizes into four discrete groups based on their relative area in the image space. We then analyze the identifiability rates of object parts across these bin groups. Figure 8 presents identifiability scores for few object-part pairs such as eye - cat, head - cow, and leg - horse, as well as aggregate trends over all parts.
@@ Appendix E Potential Application of LLaVA for Segmentation Task
For comparison, we benchmark against CLIPSeg (Lüddecke and Ecker, 2022), a model fine-tuned for segmentation and among the top-performing entries on the OVParts leaderboard. Unlike LLaVA models, CLIPSeg benefits from supervision explicitly tailored to the segmentation task, allowing it to generate high-fidelity pixel-wise masks. We report the mean IoU (mIoU) scores (Long et al., 2015) in Table 2. We find that LLaVA 7B and 13B models achieve comparable mIoU scores, but they underperform the CLIPSeg baseline considerably. The performance gap is expected given LLaVA’s lack of segmentation-specific training and the mismatch between patch-level predictions and pixel-level evaluation metrics. However, the better-than-random mIoU scores again attest to LLaVA’s ability to identify object parts and suggest that the model could be a good starting point for dedicated segmentation training.

### v2

@@ Appendix A Dataset Filtering
In order to maintain annotation consistency and visual clarity, we specifically focus on the subset of animal classes within the Pascal Panoptic Parts (Pascal-PP) dataset (de Geus et al., 2021; Meletis et al., 2020) dataset as described in Section˜3. The choice is motivated by the relatively uniform and well-defined part structures of animals, as well as their good coverage in the vocabulary of the VLMs we evaluate. To reduce noise and complexity further, we sequentially apply the following filtering criteria:
@@
After filtering, we sample up to 100 images per class, where available. Table˜1 provides further details about the distribution of images across different object classes.
@@ Appendix B Additional Results
Figures˜10, 11, 12 and 13 presents the layer-wise evolution of per-object part identifiability across different attention knockout settings described in the Section˜4.3 for TinyLLaVA 1.1B, BakLLaVA 7B, LLaVA 7B and LLaVA 13B models.
Figures˜14, 16 and 15 presents the layer-wise evolution of per-object part identifiability when progressively modifying self-attention across different layers of the CLIP Encoder for BakLLaVA 7B, LLaVA 7B and LLaVA 13B models.
@@ Appendix C Relationship between identifiability and frequency of object-parts
The heatmap in Figure˜7 presents the co-occurrence counts between different objects (columns) and their associated parts (rows), whereas the bar plot displays the overall frequency of individual parts irrespective of object association. We can observe that co-occurrence frequencies vary widely across object-part pairs. Common parts such as head, leg, and eye frequently appear with the person class, whereas parts like hoof and wing are more sparsely represented, mainly occurring with specific objects like horse and bird respectively. One might hypothesize that object parts associated with frequently occurring classes – particularly person would exhibit higher identifiability. However, analysis of OPI scores (see Table˜3) reveals no clear correlation between co-occurrence frequency and identifiability. In fact, parts with relatively low frequency often achieve high identifiability. Similarly, in Figure˜4, parts such as eye in the context of dog and cat, and leg in the context of bird and person are identified accurately despite their lower co-occurrence counts. These findings suggest that the identifiability of object parts in VLMs is not predominantly driven by training data frequency. Instead, VLMs are capable of implicitly learning object-part associations leveraging strong visual or semantic priors, and exhibit effective generalization capabilities, even without explicitly being trained for fine-grained object parts recognition.
@@ Appendix D Relationship between identifiability rate and size of object parts
In this section, we investigate whether the identifiability of object parts is influenced by their relative size. Intuitively, one might expect larger parts to be easier to identify due to their increased visibility and spatial prominence in images. To test this, we consider the CLIP model (with best-performing configuration $l=22$; see Section˜4.5) and three LLaVA variants: BakLLaVA 7B, LLaVA 7B, and LLaVA 13B. We begin by binning object parts sizes into four discrete groups based on their relative area in the image space. We then analyze the identifiability rates of object parts across these bin groups. Figure˜8 presents identifiability scores for few object-part pairs such as eye - cat, head - cow, and leg - horse, as well as aggregate trends over all parts.
@@ Appendix E Potential Application of LLaVA for Segmentation Task
For comparison, we benchmark against CLIPSeg (Lüddecke and Ecker, 2022), a model fine-tuned for segmentation and among the top-performing entries on the OVParts leaderboard. Unlike LLaVA models, CLIPSeg benefits from supervision explicitly tailored to the segmentation task, allowing it to generate high-fidelity pixel-wise masks. We report the mean IoU (mIoU) scores (Long et al., 2015) in Table˜2. We find that LLaVA 7B and 13B models achieve comparable mIoU scores, but they underperform the CLIPSeg baseline considerably. The performance gap is expected given LLaVA’s lack of segmentation-specific training and the mismatch between patch-level predictions and pixel-level evaluation metrics. However, the better-than-random mIoU scores again attest to LLaVA’s ability to identify object parts and suggest that the model could be a good starting point for dedicated segmentation training.